OpenStudy (anonymous):

Is the equation y = 5(0.5)x an example of growth or decay?

OpenStudy (anonymous):

hint: if for increasing values of x the values of y increases then its an example of growth if for increasing values of x the values of y decreases then its an example of decay hope it helps

OpenStudy (anonymous):

just think over it , its easy

OpenStudy (campbell_st):

if the power or exponent is positive its a growth model, if negative its a decay model.

OpenStudy (anonymous):

so its growth ._

OpenStudy (anonymous):

ofcourse

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

ya 100%

OpenStudy (anonymous):

how about y = 8x

OpenStudy (anonymous):

its also a growth

OpenStudy (anonymous):

how is it a decay

OpenStudy (anonymous):

its not a decay its a growth

OpenStudy (anonymous):

What is the y-intercept of the function y = 0.5(8)x? Answer 8 0.5 1

OpenStudy (campbell_st):

its linear... the power of x is always 1 for an exponential model, the power keeps increasing or decreasing 

OpenStudy (anonymous):

thanksthe help

OpenStudy (anonymous):

y- intercept = 0

OpenStudy (anonymous):

0.5

OpenStudy (anonymous):

no its zero not 0.5

OpenStudy (anonymous):

y = 200(5)x an example of growth or decay

OpenStudy (anonymous):

its a growth

OpenStudy (anonymous):

y = 1.2(3.8)x an example of growth or decay?

OpenStudy (anonymous):

growth